OpenStudy (aaliyahsims):

Estimate 5 4/10 x 7 4/5 to the nearest whole number

OpenStudy (sshayer):

\[5\frac{ 4 }{ 10 }\times7\frac{ 4 }{ 5 }=\frac{ 54 }{ 10 }\times \frac{ 39 }{ 5 }=\frac{ 1053 }{ 25 }=43.12=?\]

OpenStudy (sshayer):

or 5.4 *7.8 5.4=? write wholenumber 7.8=? write whole number. find the product.

OpenStudy (mathstudent55):

5 4/10 is just under 5.5 7 4/5 is just under 8 5 * 8 = 40 6 * 8 = 48 The average of 40 and 48 is 44. Since both numbers are just under their approximations, I'd say the answer is around 42 or 43.